OpenStudy (anonymous):

Keisha made a conjecture about this sequence: 2, 4, 8, ... Keisha’s conjecture: Since 8 is 2004-01-03-01-00-v2_files/i0020000.jpg, the pattern is to multiply the previous two elements together to get the next element. Therefore the next element would be 2004-01-03-01-00-v2_files/i0020001.jpg. Evaluate Keisha’s reasoning and create a different conjecture for this sequence.

OpenStudy (texaschic101):

Keisha is multiplying 2 * 4 to get 8......her next step would be 8 * 4 = 32. My conjecture would be your are multiplying every term by 2 to get the next term...this is a geometric sequence with a common ratio of 2 2 * 2 = 4......4 * 2 = 8.....8 x 2 = 16

OpenStudy (anonymous):

thank you